OpenStudy (anonymous):

Which of the following equations has the solution x=5

OpenStudy (aihberkhan):

Just give me the equations and I will help1 :) @Cruznatalie150

OpenStudy (alexandervonhumboldt2):

(a-b)=0 where b is the solution

OpenStudy (anonymous):

select alll that apply these are multiple choice answers

OpenStudy (alexandervonhumboldt2):

i dont see the answers

OpenStudy (anonymous):

1. 2^3x=8^3x

OpenStudy (anonymous):

2. 4^x=4^5x-20

OpenStudy (alexandervonhumboldt2):

just substitute 5 for x in each equation and check which will be correct

OpenStudy (anonymous):

3. 4^x+6=9^x+8

OpenStudy (anonymous):

4. 2^x=6^x-3

OpenStudy (anonymous):

i did but i got stuck on 3

OpenStudy (anonymous):

this is simplifying radicals

OpenStudy (alexandervonhumboldt2):

4^3==4*4*4=64

OpenStudy (anonymous):

if you could explain to me step by step

OpenStudy (anonymous):

but its just 4^x it dosent have a varriable

OpenStudy (welshfella):

when x = 5, 4^x = 4^5 = 1024

OpenStudy (welshfella):

4*4*4*4*4 = 4^5

OpenStudy (anonymous):

oh ok so then what would be the answer or how would i solve it @welshfella

OpenStudy (anonymous):

select alll that apply

OpenStudy (welshfella):

for number 2 left side = 4^5 = 1024 right side = 4^(5(5) - 20) = 4^5 = 1024 so?

OpenStudy (anonymous):

@AlexandervonHumboldt2 @welshfella like for is equal to 5 right question 4 i substituted them in divided by 2 both sides

OpenStudy (anonymous):

beause it asking has a solution to 5

OpenStudy (anonymous):

so then 2 would work

OpenStudy (welshfella):

?? the left side = right side Therefore x = 5 is a solution to this equation.

OpenStudy (welshfella):

Yes equation 2 has the solution x = 5.

OpenStudy (anonymous):

DOES EQUATION 4 HAVE SOLUTION FOR 5 I THINK IT DOES BUT AM I RIGHT

OpenStudy (welshfella):

No LHS = 2^5 = 32 RHS = 6^(5-3) = 6^2 = 36